OpenStudy (anonymous):

Use the functions m(x) = 5x + 4 and n(x) = 6x − 9 to complete the function operations listed below. Part A: Find (m + n)(x). Show your work. (3 points) Part B: Find (m ⋅ n)(x). Show your work. (3 points) Part C: Find m[n(x)]. Show your work. (4 points)

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (e.mccormick):

Well, do you know the process in general for this sort of thing?

OpenStudy (kash_thesmartguy):

Maybe, this will help @kkailyn - http://www.purplemath.com/modules/fcnops.htm

OpenStudy (study_together):

\[m=5x+4\]\[n=6x-9\]Add both of those together to get your answer. Multiply both of them to get your answer. Distribute both of them to get your answer.

OpenStudy (study_together):

Can I get a medal?

OpenStudy (anonymous):

well yes i do i just dont know how to word it @e.mccormick and @Study_together why?

OpenStudy (e.mccormick):

Well, function composition is mainly an exercise in writing things out properly, then simplifying. If I said simplify x + 2 + x + 5 it is pretty obvious it is 2x + 7. But if I say f(x)= x + 2 and g(x) = x + 5 then ask "What is (f + g)(x)?" it is less obvious, but it is the same thing. It just means you write out some more steps: (f + g)(x) = f(x) + g(x) (f + g)(x) = (x + 2) + (x + 5) (f + g)(x) = x + 2 + x + 5 (f + g)(x) = 2x + 7 That shows the steps, basically. Do similar things with your functions.

OpenStudy (anonymous):

for part b do i just multiply the functions instead of adding like i did for the 1st one?

OpenStudy (e.mccormick):

b is multiply. c is function inside a function. So you replace x in the outter function with the entore inner. So to take my example, f[g(x)] means I do this: x+2 (x+5)+2 x+7 I am putting the x+5 inside the x+2 in place of the x. So again, just use your functions and do the same thing.